OpenStudy (anonymous):

Given the two functions below, what is the y-intercept of g(x)? f(x) = x2 g(x) = f(x –7) Type your answer as an integer, like this: -14

OpenStudy (johnweldon1993):

To find the y-intercept...we would make x = 0 g(x) = f(x - 7) where we see that 'x' we would replace it with whatever f(x) equals...so g(x) = x² - 7 If we want the y-intercept...x must = 0 .....so we have g(x) = 0 - 7 the y-intercept must = -7 If i'm correct

jimthompson5910 (jim_thompson5910):

you replace x with (x-7), so g(x) = (x-7)^2

OpenStudy (anonymous):

Can I have your brains John and Jim? Lol

OpenStudy (anonymous):

jim's answer is correct, the answer is not -7

OpenStudy (johnweldon1993):

Wait @jim_thompson5910 did I miss that? hang on lol

jimthompson5910 (jim_thompson5910):

lol no sorry, but I'm sure you have brains as well brandeejo

OpenStudy (anonymous):

it should be 49

OpenStudy (anonymous):

(0-7)^2 = 49

jimthompson5910 (jim_thompson5910):

so all you do is plug in x = 0 to get... g(x) = (x-7)^2 g(0) = (0-7)^2 g(0) = (-7)^2 g(0) = 49

OpenStudy (anonymous):

Im only kidding. thank you guys. All have been such a huge help!

jimthompson5910 (jim_thompson5910):

so you know the point (0,49) is on g(x)

OpenStudy (johnweldon1993):

AHH! thank you sorry blind moment there lol

jimthompson5910 (jim_thompson5910):

that's fine

OpenStudy (johnweldon1993):

So yeah ignore my post @brandeejo :)

OpenStudy (anonymous):

Your fine John!